step by step plz

Example. For the given matrix below compute both det(A) and det (5.A). 025 A=|1 7 10 01 3
Answer #1

To 2 51 col:- Given A = lia 10 lo i 3) +5 det A = 017 101-2 o 3 = 0 (21-10) - 2 (3-0) +5(1-0) = 0 - 6+5 deta = -1 To 2 57 Now